OpenStudy (anonymous):

Let f(x)=1/x. Compute the difference quotient for f(x) at x=5 with h=0.2 Use the Power Rule to compute the derivative: d/dt, t2/3|t=6

OpenStudy (kirbykirby):

\[\large \frac{f(x+h)-f(x)}{h} = \frac{\frac{1}{x+h}-\frac{1}{x}}{h}\] Now you just need to simply substitute \(x=5\) and \(h=0.2\) into the above which should be straightforward (I hope)!

OpenStudy (anonymous):

okay, so 1/5.2 - 1/5 = .2??

OpenStudy (anonymous):

so would the derivative be 1?

OpenStudy (kirbykirby):

Well more like 1/5.2 - 1/5 , and then divide that whole result by 0.2

OpenStudy (kirbykirby):

hm that subtraction shouldn't give 0.2 o_O

OpenStudy (anonymous):

what are you getting?

OpenStudy (kirbykirby):

\(\dfrac{1}{5.2}-\dfrac{1}{5}=-0.007692\)

OpenStudy (anonymous):

yeah that's what i got on my calculator, so i do -.007692/.2?

OpenStudy (kirbykirby):

yep \(\checkmark\)

OpenStudy (anonymous):

i'm getting -.384615385

OpenStudy (anonymous):

-.0384615385**

OpenStudy (kirbykirby):

yup !

OpenStudy (anonymous):

how about the second part of the problem?

OpenStudy (anonymous):

Use the Power Rule to compute the derivative: d/dt, t2/3|t=6

OpenStudy (kirbykirby):

I just want to be clear since the notation doesn't translate well in regular text... the problem is asking for: \[ \frac{d}{dt}\, \left.\frac{t^2}{3}\right|_{t=6} \] ?

OpenStudy (anonymous):

that's correct, except it should be t^(2/3)

OpenStudy (kirbykirby):

ohh ok I see

OpenStudy (kirbykirby):

So recall that the power rule states: \[\frac{d}{dt} t^n=nt^{n-1} \], here your \(n=\dfrac{2}{3}\)... so: \[\large \frac{d}{dt} t^{\frac{2}{3}}=\frac{2}{3}t^{\frac{2}{3}-1}=\frac{2}{3}t^{-\frac{1}{3}}\]

OpenStudy (kirbykirby):

Now, since they want the derivative evaluated at t=6, just substitute t=6 into the above

OpenStudy (anonymous):

so 0.366881?

OpenStudy (kirbykirby):

ya that would be good :) (But if your question requires an exact answer rather than a decimal approximation... then you wouldn't need to evaluate it further)

OpenStudy (anonymous):

thank u!

OpenStudy (kirbykirby):

your welcome =]